OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? Answer 27 54 62 81

OpenStudy (anonymous):

b + g = 89 b = 2g + 8 => b = 2(89-b) + 8 => 3b = 2*89 + 8 => 3b = 186 => b = 186/3 = 62